# Fully Energy-Efficient Randomized Backoff: Slow Feedback Loops Yield Fast Contention Resolution

### v2

@@ 1.1. Model
Finally, we observe that there are several natural settings in which implicit-throughput guarantees directly imply strong guarantees on packet backlog, even for infinite input sequences. Suppose, in particular, that packets arrive according to adversarial queuing theory, which parameterizes the “burstiness” of packet-arrival in infinite streams. In adversarial queuing theory (BenderFaHe05, ; ChlebusKoRo12, ; BKS14, ; ChlebusKoRo06, ; BorodinKlRaSu01, ; BorodinKlRa96, ), the adversary is restricted from injecting too many packets/jammed slots over a set of consecutive slots of length $S$, where $S$ is a parameter of the model that we refer to as the granularity. For granularity $S$, the number of packet arrivals plus the number of jammed slots is limited to $\lambda S$, where the arrival rate $\lambda$ is a constant less than one. On the other hand, how the packet arrivals are distributed within each $S$-sized window is adversarial: no restrictions are placed on how the (at most) $\lambda S$ packets and jammed slots are distributed. By showing that the implicit throughput of all active slots is $\Omega(1)$, we obtain as a corollary a strong bound of $O(S)$ on the number of packets backlogged in the system at all times, as long as $\lambda$ is a sufficiently small constant.
1.2. Main results
@@
Our algorithm relies on the natural multiplicative-weight approach to backoff—with a careful choice of probabilities and updates rules. In contrast to some previous approaches, we do not rely on packet batching (to turn the online problem into a series of batch problems), leader election, busy tones, population estimation, dividing the channel into simulated subchannels (e.g., odd and even slots), or other approaches seen in many modern algorithms; alas, these seem problematic in our setting (and potentially complicated to implement). Given the simplicity of our algorithm, the technical innovation lies in choosing the update parameters, analyzing the underlying combinatorial process, and proving that it is fast, robust, and fully energy efficient.
@@
Theorem 3 (Implicit throughput, infinite packet streams).
@@
Corollary 4 (Throughput, finite packet streams).
@@
Corollary 5 (Bounded backlog for adversarial-queuing arrivals).
@@
Theorem 6 (Energy, finite executions).
@@
Theorem 7 (Energy, adversarial queuing).
@@
Theorem 8 (Energy, infinite executions).
@@
Theorem 3 is proved in Section 5.5 where it appears as Corollary 11. Corollary 5 corresponds to Corollary 14. The remaining theorems are proved in Section 5.6. In particular, Theorem 6 corresponds to Theorem 3.4, and Theorem 7 corresponds to Theorem 16. Finally, Theorem 8 is included in Theorem 3.4.
1.3. Extensions to Reactive Adversary
A reactive adversary (richa:efficient-j, ; wilhelm2011short, ; DBLP:journals/dc/KingPSY18, ) has an instantaneous reaction time; that is, this adversary listens to the channel and can decide whether to jam and/or inject new packet(s) in slot $t$ based on what it hears in slot $t$ itself. In contrast, the standard adaptive adversary would not know whether any packet chooses to send in slot $t$ until slot $t+1$. This allows a reactive adversary to cheaply prevent any particular packet $p$ from transmitting successfully by jamming only those slots where $p$ makes transmission attempts. Consequently, against a reactive adversary, the total number of channel accesses required is at least linear in the amount of jamming. In the case of exponential backoff, the situation is more dire: for any $T$ a reactive adversary can also drive the throughput down to $O(1/T)$ by jamming a single packet a mere $\Theta(\log T)$ times.
Note that reactivity and adaptivity are somewhat orthogonal. Reactivity addresses how quickly the adversary can react to the visible state on the channel—importantly, only sending is revealed in this way, since it is visible. In contrast, the adaptive adversary knows all of the internal state and random choices of packets up to the previous slot, and in particular this adversary also knows if packets choose to listen.
In addition to our main results on purely adaptive adversaries, this paper also addresses an adversary that is both adaptive and reactive. It turns out that the reactive adversary does not impact the implicit-throughput bounds for our algorithm (our analysis applies whether or not the adversary can see the activity on the channel at the current time). Reactivity thus only impacts the number of channel accesses. Roughly speaking, the theorem states that the reactive adversary has nontrivial impact on the worst-case number of channel accesses and thus energy, which is to be expected, but it does not have significant impact on the average.
Theorem 9 (Energy, reactive adversary).
@@ 1.3. Extensions to Reactive Adversary
The items of Theorem 9 are each proved separately as Theorems 15, 17, and 3.4 in Section 5.6.
2. Low-Sensing Backoff Algorithm
This section presents the Low-Sensing Backoff algorithm; see Figure 1.
@@ 2. Low-Sensing Backoff Algorithm
For ease of presentation, we describe our algorithm as listening whenever it sends. However, the packet need not actually do both; observe that any packet that is sending does not need to listen to determine the state of the channel, since if the packet is still in the system after sending in slot $t$, then slot $t$ was noisy.

### v5

@@ 1.1 Model
Finally, we observe that there are several natural settings in which implicit-throughput guarantees directly imply strong guarantees on packet backlog, even for infinite input sequences. Suppose, in particular, that packets arrive according to the adversarial queuing theory arrival model (described below), which parameterizes the “burstiness” of packet-arrival in infinite streams. In adversarial queuing theory [BenderFaHe05, ChlebusKoRo12, ChlebusKoRo06, BorodinKlRaSu01], the adversary is restricted from injecting too many packets/jammed slots over a set of consecutive slots of length $S$, where $S$ is a parameter of the model that we refer to as the granularity. For granularity $S$, the number of packet arrivals plus the number of jammed slots is limited to $\lambda S$, where the arrival rate $\lambda$ is a constant less than one. On the other hand, how the packet arrivals are distributed within each $S$-sized window is adversarial: no restrictions are placed on how the (at most) $\lambda S$ packets and jammed slots are distributed. By showing that the implicit throughput of all active slots is $\Omega(1)$, we obtain as a corollary a strong bound of $O(S)$ on the number of packets backlogged in the system at all times, as long as $\lambda$ is a sufficiently small constant.
1.2 Main results
@@
Our algorithm relies on the natural multiplicative-weight approach to backoff—with a careful choice of probabilities and update rules. In contrast to some previous approaches, we do not rely on packet batching (to turn the online problem into a series of batch problems), leader election, busy tones, population estimation, dividing the channel into simulated subchannels (e.g., odd and even slots) for packet coordination, or other approaches seen in many modern algorithms; alas, these seem problematic in our setting. Given the simplicity of our algorithm, the technical innovation lies in choosing the update parameters, analyzing the underlying combinatorial process, and proving that it is fast, robust, and fully energy-efficient.
@@
Theorem 1.1 (Implicit throughput, infinite packet streams).
@@
Corollary 1.2 (Throughput, finite packet streams).
@@
Corollary 1.3 (Bounded backlog for adversarial-queuing arrivals).
@@
Theorem 1.4 (Energy, finite executions).
@@
Theorem 1.5 (Energy, adversarial queuing).
@@
Theorem 1.6 (Energy, infinite executions).
@@
The above statements are simplified slightly for ease of presentation, while the formal statements are presented later in Section LABEL:sec:analysis. Specifically, Theorem 1.1 is proved in Section LABEL:sec:bettinggame where it appears as Corollary LABEL:cor:implicit-throughput, and Corollary 1.3 corresponds to Corollary LABEL:cor:backlog. The remaining theorems are proved in Section LABEL:sec:energy. In particular, Theorem 1.4 corresponds to Theorem LABEL:thm:finite_adaptive_accesses, and Theorem 1.5 corresponds to Theorem LABEL:thm:queuing_adversarial_accesses. Finally, Theorem 1.6 is included in Theorem LABEL:thm:infinite_accesses.
1.3 Extensions to Reactive Adversary
A
Definition 1.
reactive adversary [richa:jamming3, wilhelm2011short, DBLP:journals/dc/KingPSY18] has an instantaneous reaction time; that is, this adversary listens to the channel and can decide whether to jam and/or inject new packet(s) in slot $t$ based on what it hears in slot $t$ itself. In contrast, the standard adaptive adversary would not know whether any packet chooses to send in slot $t$ until slot $t+1$. This allows a reactive adversary to cheaply prevent any particular packet $p$ from succeeding by jamming only those slots where $p$ makes transmission attempts. Thus, against a reactive adversary, the total number of channel accesses required is at least linear in the amount of jamming. For exponential backoff, the situation is more dire: for any $T$ a reactive adversary can also drive the throughput down to $O(1/T)$ by jamming a single packet a mere $\Theta(\ln T)$ times.
Note that reactivity and adaptivity are somewhat orthogonal. Reactivity addresses how quickly the adversary can react to the detectable channel state—importantly, only sending is revealed, since it is detectable. In contrast, an adaptive adversary knows all of the internal state and random choices of packets up to the previous slot, and in particular this adversary also knows if packets choose to listen.
In addition to our main results on purely adaptive adversaries, we also address an adversary that is both adaptive and reactive. It turns out that the reactive adversary does not impact our implicit-throughput bounds for our algorithm (our analysis applies whether or not the adversary can see the channel activity at the current time). Reactivity thus only impacts the number of channel accesses. Roughly speaking, the theorem states that the reactive adversary has nontrivial impact on the worst-case number of channel accesses and thus energy, which is to be expected, but it does not have significant impact on the average.
Theorem 1.7 (Energy, reactive adversary).
@@ 1.3 Extensions to Reactive Adversary
The items of Theorem 1.7 are each proved separately as Theorems LABEL:thm:finite_reactive_accesses, LABEL:thm:queuing_reactive_accesses, and LABEL:thm:infinite_accesses in Section LABEL:sec:energy.
2 Related Work
Contention resolution has been an active area of research for several decades. Here, we discuss related results from the literature. We highlight that a preliminary version of this work has appeared previously [bender2024fully].
Robustness.
Early work on jamming resistance by Awerbuch, Richa, and Scheideler [awerbuch:jamming] considers the case of $N$ devices, each of which always has a packet to ready to send; that is, rather than changing over time, the number of packets contending for channel access is always $N$. For any interval of at least $T$ consecutive slots, at most a $(1-\lambda)$-fraction may be jammed, where the constant $\lambda>0$ is unknown. In this setting, Awerbuch, Richa, and Scheideler[awerbuch:jamming] showed that constant throughput is possible with a polylogarithmic number of transmissions per packet. Several subsequent results have built on this result, addressing multi-hop [richa:jamming2, richa:competitive-j] and co-located networks [richa:jamming4], reactive jamming [richa:jamming3, richa:efficient-j], and aspects of signal propagation [ogierman:competitive]. However, this series of results relies on devices knowing rough bounds on $N$ and $T$; in particular, devices must set a parameter whose value is $O(1/(\ln\ln(N)+\ln(T)))$.
A similar restriction on jamming is explored by Anantharamu et al. [anantharamu2019packet, anantharamu2011medium], where $\rho t+b$ slots in any window of consecutive $t$ slots can be jammed, for $0<\rho\leq 1$, and a non-negative integer, $b$. Probabilistic jamming has also been considered by Chlebus et al. [Chlebus:2016:SWM:2882263.2882514] in a multi-channel setting, and by Agrawal et al. [DBLP:conf/spaa/AgrawalBFGY20] where packets have delivery deadlines.
When the amount of jamming can be unbounded, Bender et al. [BenderFiGi19, BenderFiGi16] showed expected constant throughput is possible over non-jammed slots using a $\operatorname{polylog}(N)$ number of sending attempts per packet. Chang, Jin, and Pettie [ChangJiPe19] give an elegant algorithm that achieves the same asymptotic result, but drastically improves the expected throughput, rivaling the $1/e$ result achieved by the well-known slotted ALOHA [binder:aloha], despite the more-challenging adversarial setting.
There is a growing body of work [ChenJiZh21, jiang:robust, DeMarcoSt17, BenderKoKu20] on how to design contention resolution protocols where packets listening on the channel during a given slot learn whether or not that slot was successful, but not whether the slot was empty or noisy; that is, packets lack the ability to detect collisions. As mentioned earlier, the results by De Marco and Stachowiak [DeMarcoSt17], De Marco, Kowalski, and Stachowiak [marco:time], and Bender et al. [BenderKoKu20] address this setting. Dealing with jamming in the absence of collision detection is challenging. Here, results by Chen, Jiang, and Zheng [ChenJiZh21] and Jiang and Zheng [jiang:robust] show how jamming can still be tolerated, although throughput is $\Theta(1/\ln N)$ and packets are allowed to listen on the channel for free.
There is a large body of work on applied security considerations with respect to jamming. The impact of these attacks has been evaluated under a variety of wireless settings and adversarial strategies [bayraktaroglu:performance, xu:feasibility, lin:link, aschenbruck:simulative], and several defenses have been proposed. For example, spread-spectrum techniques involve devices coordinating their communication over different portions of the spectrum in order to evade collisions [spread:foiling, navda:using, deejam:wood]. Another defense is to have devices map the areas impacted by jamming and then routing traffic around those areas [wood2003jam, han2018hierarchical]. Specialized antenna technology can be used to filter out a jamming signal, which is a process referred to as null steering [zheng2018design, mcmilin2015gps].
Energy efficiency.
As discussed in Section 1, there are a handful of results on contention resolution that achieve constant throughput and address full energy efficiency. To reiterate, Bender et al. [BenderKoPe18] achieve expected $O(\ln(\ln^{*}N))$ channel accesses per packet, which Chang et al. [ChangKoPe19] prove is tight. De Marco and Stachowiak [DeMarcoSt17] and De Marco, Kowalski, and Stachowiak [marco:time] give sending-efficient algorithms and conjecture that it is possible to extend the results in their papers so that each packet listens to the channel a polylogarithmic number of times (see Section 7 Discussion and open problems in [marco:time]). However, these results deviate significantly from the ternary model (requiring control messages and, in the case of [DeMarcoSt17, marco:time], Good Samaritan packets).
We also note that full energy efficiency has drawn attention outside the domain of contention resolution. Many of these results address wireless networks where, in order to extend the network’s operational lifetime, low-power devices may spend significant time in an energy-efficient sleep state, waking up rarely to access the channel (e.g., [wang2019connectivity, deng2005scheduling, li2005energy, ye2006ultra, ullah2009study, merlin2010duty]). Several algorithmic results address energy efficiency in the context of traditional distributed computing problems(e.g., [BarenboimM21, chatterjee2020sleeping, ghaffari2023distributed, dufoulon2023distributed, chang2018energy, king:sleeping, dani2021wake]).
Dynamic packet arrivals.
Much of the initial work on contention resolution addresses arrival times that are dictated by a stochastic process [MetcalfeBo76, Goldberg-notes-2000, hastad:analysis, goldberg2004bound, shoch1980measured, abramson1977throughput, kelly1987number, aldous:ultimate, GoodmanGrMaMa88, mosely1985class, RaghavanU98, RaghavanU95, GOLDBERG1999232]. The intricacies of whether the contention-resolution algorithm knows local or global history, can adapt its actions over time, can listen on the channel, knows the number of packets in the system, etc. have been explored in depth [tsybakov1979ergodicity, ferguson1975control, fayolle1977stability, rosenkrantz1983instability, kelly:decentralized, paterson1995contention, DBLP:journals/jacm/GoldbergMPS00, tsybakov1978free-russian, tsybakov1978free, hayes:adaptive, massey1981collision, gallager, tsybakov1980random, verdu1986computation, tsybakov1987upper]. A survey of much of the literature in this setting is provided by Chlebus [chlebus2001randomized].
In contrast to stochastic arrivals, packets may arrive at times scheduled by an adversary. The special case where all packets arrive simultaneously has been explored, and $O(n)$ makespan is possible [Gereb-GrausT92, GreenbergL85, FernandezAntaMoMu13, BenderFaHe05, BenderFiGi06]. More generally, when packets are injected over time in a worst-case fashion, models from adversarial queuing theory [BorodinKlRa96, cruz-one, andrews2001universal] have been used to derive stability results [chlebus2009maximum, chlebus2007stability, anantharamu2010deterministic, aldawsari:adversarial, anantharamu2019packet, anantharamu2011medium, anantharamu2015broadcasting, hradovich2020contention, bienkowski:distributed, garncarek2018local, garncarek2019stable]. Several results explore deterministic approaches for contention resolution (e.g., [deMarco:fast, greenberg1985lower, DBLP:conf/icdcs/MarcoKS19, DEMARCO20171, de2013contention, anantharamu2009adversarial, anantharamu2017adversarial, Kowalski05], although such approaches often have bounds that are asymptotically inferior to their corresponding randomized solutions.
Other computation on a shared channel.
A closely-related problem to contention resolution is the wake-up problem, which addresses how long it takes until a single successful transmission [jurdzinski2015cost, jurdzinski2005probabilistic, jurdzinski2002probabilistic, jurdzinski2002probabilistic, farach2007initializing, farach2015initializing, Chlebus:2016:SWM:2882263.2882514, ChrobakGK07, ChlebusGKR05, DBLP:journals/dam/MarcoPS07, DBLP:conf/ictcs/MarcoPS05, willard:loglog, fineman:contention, fineman:contention2]; this is in contrast to having all packets succeed.
More generally, several fundamental distributed computing challenges have been studied involving communication on a multiple access channel, and often even in the presence of adversarial jamming. These include broadcast [koo2, gilbert:near, gilbert:making, chen:broadcasting], leader election [gilbert:malicious, gilbert2009malicious, DBLP:journals/talg/AwerbuchRSSZ14], gossip [gilbert:interference, dolev:gossiping], node discovery [meier:speed], and simple point-to-point communication [king:conflict, DBLP:journals/dc/KingPSY18], packet scheduling [DBLP:conf/waoa/JurdzinskiKL14, DBLP:journals/tcs/AntaGKZ17].
3 Low-Sensing Backoff Algorithm
@@
This section presents the Low-Sensing Backoff algorithm; see Figure 1. For ease of presentation, we describe our algorithm as listening whenever it sends. However, the packet need not actually do both; observe that any packet that is sending does not need to listen to determine the state of the channel, since if the packet is still in the system after sending in slot $t$, then slot $t$ was noisy.
Low-Sensing Backoff for packet $u$ Key Variables: $w_{u}(t)$: window size of $u$ in slot $t$.
If $u$ is injected at time slot $t$, then $w_{u}(t)=w_{\rm min}$.
$c$: a sufficiently large positive constant.